OpenStudy (anonymous):

using chain rule: y=5cos^-4x

OpenStudy (anonymous):

You want to find derivative here ???

OpenStudy (anonymous):

yes please.

OpenStudy (anonymous):

Firstly tell me what is the derivative of cos(x) ???

OpenStudy (anonymous):

-sinx

OpenStudy (anonymous):

Yep.. Your question is : \[y = 5\cos^{-4}(x)\] Right ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know how to find derivative when power is there?? Here power is -4 ..

OpenStudy (anonymous):

\[\frac{d}{dx}(f^n) = n \cdot f^{n-1}\]

OpenStudy (anonymous):

yes I know that

OpenStudy (anonymous):

\[\frac{d}{dx}(\cos^{-4}(x)) = -4 \cdot \cos^{-5}(x) \cdot \frac{d}{dx}(\cos(x)) \cdot \frac{d}{dx}(x)\]

OpenStudy (anonymous):

Can you go forward from here ???

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

If problem arises, then @UnkleRhaukus and @sauravshakya will help you in going forward.. Gotta go...